# Semi-empirical constraints on the HI mass function of star-forming galaxies and $\Omega_{\rm HI}$ at $z\sim 0.37$ from interferometric surveys

### v1

@@ Abstract
Results. We find evolution of the HIMF at $z\sim 0.37$ with respect to results at $z\sim 0$ from the ALFALFA survey and at $z\sim 1$ from uGMRT data. Our results for $\Omega_{\rm HI}$ are in broad agreement with other literature results, and follow the overall trend on $\Omega_{\rm HI}$ as a function of redshift. The derived value $\Omega_{\rm HI}=\left(7.02^{+0.59}_{-0.52}\right)\times 10^{-4}$ at $z\sim 0.37$ from the combined analysis deviates at $\sim 2.9\sigma$ from the ALFALFA result at $z\sim 0$.
Conclusions. We conclude that the HIMF and $\Omega_{\rm HI}$ constraints that we derive from state-of-the-art deep interferometric surveys suggest an evolution of the HIMF and of the cosmic H I density, supporting a picture of smooth transition of the H I content of galaxies from $z\sim 0$ to $z\sim 1$.
@@ 1 Introduction
The neutral atomic hydrogen (hereafter H I) plays a key role in astrophysics and cosmology. In fact, the distribution of H I in the Universe is driven by the complex interplay of astrophysical and cosmological effects. The abundance of H I in the large-scale structure is regulated by the nonlinear formation and evolution of the cosmic web (e.g., Martizzi et al., 2019; Galárraga-Espinosa et al., 2021; Sinigaglia et al., 2021, 2024). Once galaxies form within dark matter haloes thanks to the conversion of the intra-halo H I to molecular hydrogen and to gravitational collapse (e.g., Somerville & Davé, 2015; Vogelsberger et al., 2020; Primack, 2024), the H I constitutes the main raw fuel for star formation and is one of the major ingredients contributing to the baryon cycle and in regulating the galaxy mass assembly (e.g., Blitz & Rosolowsky, 2006; Bigiel et al., 2008; Sternberg et al., 2014). For these reasons, constraining the H I cosmic density parameter $\Omega_{\rm HI}$ and the H I mass function (hereafter HIMF) over cosmic time is of paramount importance to try answering profound questions about the Universe, such as probing cosmic reionization (e.g., Koopmans et al., 2015; DeBoer et al., 2017), understanding the galaxy mass assembly and the time evolution of the star formation rate density (e.g., Huang et al., 2012; Maddox et al., 2015; Catinella et al., 2018; Sinigaglia et al., 2022; Bianchetti et al., 2025), and even using H I to infer cosmological information such as e.g. unveiling the nature of dark matter (e.g., Pinetti et al., 2020; Bauer et al., 2021; Bañares-Hernández et al., 2023) and of dark energy (e.g., Santos et al., 2015; Zhang et al., 2023; Berti et al., 2024), as well as testing the $\Lambda$CDM model and other alternative theories of gravity with resolved galaxy rotation curves (e.g., O’Brien et al., 2018).
@@
In this work, we follow a methodology similar in spirit to the one by Bera et al. (2022) and Chowdhury et al. (2024). In particular, we combine the most recent measurements of the stellar mass function of star-forming galaxies in the COSMOS field at $0.50.2<z<$ 0.5 by Weaver et al. (2023) and of the $M_{\rm HI}-M_{\star}$ scaling relation of star-forming galaxies at $0.490.22<z<$ 0.49 (median redshift $z\sim 0.37$) by Bianchetti et al. (2025) from the H I stacking analysis of two state-of-the-art deep H I interferometric surveys MIGHTEE (Jarvis et al., 2016) and CHILES (Fernández et al., 2013), both covering the COSMOS field (Scoville et al., 2007). With these tools, we derive novel indirect semi-empirical constraints on the HIMF of star-forming galaxies and the cosmic H I density parameters $\Omega_{\rm HI}$ at $z\sim 0.37$.
@@
Throughout this work, we adopt the following cosmological parameters: $\Omega_{m}=0.3$, $\Omega_{\Lambda}=0.7$, $H_{0}=70.0\leavevmode\nobreak\ {\rm km}\leavevmode\nobreak\ {\rm s}^{-1}\leavevmode\nobreak\ {\rm Mpc}^{-}1$. When relevant, we have rescaled the results from other literature results to the match the cosmological parameters adopted herein.
@@ 2.1 The COSMOS2020 sample and the stellar mass function
The optical catalog underlying the stacking analysis performed by Bianchetti et al. (2025) was built through a cross-match between the COSMOS2020-CLASSIC [^1] Release 1, v2.2, March 2023 photometric catalog (Weaver et al., 2022) and a spectroscopic catalog assembled by merging three different catalogs covering the COSMOS field: the COSMOS spec-$z$ compilation (Khostovan et al., 2025), the DEVILS survey catalog (Davies et al., 2018) and the DESI survey Early Data Release catalog (DESI Collaboration et al., 2023). First, Bianchetti et al. (2025) selected only spectroscopic sources in the COSMOS field (Scoville et al., 2007) and at $0.490.22<z<$ 0.49 (median redshift $z\sim 0.37$), i.e. within a subset of the volume probed by the MIGHTEE and CHILES data sets (see Sect. 2.2). In particular, they performed a positional cross-match (with matching radius=$1$ arcsec) between the spectroscopic and the photometric catalog. The photometric catalog includes galaxy properties obtained through spectral energy density (SED) fitting performed with the LePhare code (Ilbert et al., 2006; Arnouts & Ilbert, 2011). Therefore, after the cross-match, they retrieved from the photometric catalog all the aforementioned photometrically-derived parameters for each of the sources in the spectroscopic catalog. Star-forming galaxies were then selected using the in-built color-color selection in the $(NUV-r)/(r-J)$ rest-frame plane. Further quality cuts were then applied to ensure the highest possible reliability, including redshift quality cuts and photometric redshift outliers exclusion.
@@
In this work, we also make use of the results on the stellar mass function in the COSMOS field as measured from the COSMOS2020 catalog (Weaver et al., 2023). In particular, Weaver et al. (2023) performed a detailed study of the stellar mass sample from COSMOS2020, deriving the stellar mass functions for the total, star-forming, quiescent populations at $0.50.2<z<$ 0.5, as well as in several higher redshift ranges. The splitting of the galaxy population into star-forming and quiescent was performed according to the same color criterion mentioned above. These aspects, together with the fact that the underlying photometric catalog is the same, make the stellar mass function derived by Weaver et al. (2023) as the ideal candidate to be used in this work. Weaver et al. (2023) used a Markov Chain Monte Carlo method to fit a double Schechter profile to the resulting stellar mass function:
$\Phi\leavevmode\nobreak\ d\log_{10}(M_{\star})$ $=\log(10)\exp\left(-10^{\log_{10}(M_{\star})-\log_{10}(\tilde{M})}\right)\times$ (1)
×$\times\biggl{[}\tilde{\Phi}_{1}\left(10^{\log_{10}(M_{\star})-\log_{10}(\tilde{M})}\right)^{\alpha_{1}+1}+$ (2)
$+\tilde{\Phi}_{2}\left(10^{\log_{10}(M_{\star})-\log_{10}(\tilde{M})}\right)^{\alpha_{2}+1}\biggr{]}\leavevmode\nobreak\ d\log_{10}(M_{\star})\quad.$ (3)
For the star-forming mass-complete galaxy subsample at $0.50.2<z<$ 0.5, Weaver et al. (2023) found the following best-fitting parameters:
[…]
@@ 2.2 The MIGHTEE and CHILES surveys
MIGHTEE (Jarvis et al., 2016) is a survey conducted with MeerKAT radio telescope. The MeerKAT radio interferometer (Jonas & MeerKAT Team, 2016) consists of 64 offset Gregorian dishes equipped with receivers in UHF–band ($580$ MHz $<\nu<$ $1015$ MHz), L–band (900 MHz $<\nu<$1670 MHz), and S–band (1750 MHz$<\nu<$ 3500 MHz), and serves as a precursor to the Square Kilometre Array (SKA). MIGHTEE is an L-band continuum, polarization, and spectral-line large survey conducted with MeerKAT, utilizing spectral and full Stokes mode observations. It covers four deep extragalactic fields (COSMOS, XMM-LSS, ECDFS, ELAIS-S1), chosen because of their extensive multi-wavelength coverage from previous and ongoing observations. For this paper, we use results derived for the MIGHTEE-H I Early Science spectral line data from MIGHTEE (Maddox et al., 2021) covering the COSMOS field with a single pointing within the redshift range of $0.490.22<z<$ 0.49, with an effective exposure time of approximately 23 hours.
CHILES is a deep-field H I survey, carried out with the VLA and imaging H I over a contiguous redshift range of $0.490<z<$ 0.49 for the first time (Fernández et al., 2013). The pointing was centered on RA(J2000) 10h01m24s and DEC(J2000) 2d21m00s. It samples a subregion of the COSMOS field; therefore, it provides an independent measurement of the same patch of sky as the MIGHTEE survey. The 25-meter diameter antennas of the VLA provide a field of view (FOV) of around $0.5$ deg at $1.4$ GHz. Due to the rotating VLA configurations, the CHILES observations were split into five observing epochs spaced approximately 15 months apart. In this paper, we use results from successfully processed data from all five of the observing epochs, amounting to approximately $800$ observation hours (approximately $600$ on-source hours).
@@ 2.3 Summary of stacking results from MIGHTEE and CHILES
$\log_{10}(M_{\star})$ bin Median $\log_{10}(M_{\star})$ $\braket{\log_{10}(M_{\rm HI})}$
@@
$9.58.0<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 9.5 $9.10$ $9.51\pm 0.05$
$9.89.5<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 9.8 $9.65$ $9.68\pm 0.06$
$10.59.8<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 10.5 $10.17$ $9.94\pm 0.02$
$\log_{10}(M_{\star}/{\rm M_{\odot}})>10.5$ $10.85$ $10.07\pm 0.04$
@@
$9.58.0<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 9.5 $9.10$ $9.60\pm 0.04$
$10.59.5<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 10.5 $10.06$ $9.94\pm 0.02$
$\log_{10}(M_{\star}/{\rm M_{\odot}})>10.5$ $10.85$ $10.12\pm 0.03$
@@
$9.58.0<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 9.5 $9.10$ $9.55\pm 0.07$
$10.59.5<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 10.5 $10.03$ $9.82\pm 0.06$
$\log_{10}(M_{\star}/{\rm M_{\odot}})>10.5$ $10.83$ $10.13\pm 0.08$
Table 1: Data points from the H I spectral stacking analysis performed by Bianchetti et al. (2025). The left column reports the $M_{\star}$ interval underlying each of the measured data point, the central column the corresponding median $M_{\star}$ and the right column the average $M_{\rm HI}$ measured from stacking.
@@ 3.1 SIMBA
In this work, we make use of the publicly-available galaxy catalogs at $z\sim 0.365$ for SIMBA-100 and for all the SIMBA-50 boxes. Consistently with the observational sample selection by Bianchetti et al. (2025), we select only star-forming galaxies with $\log_{10}(M_{*})>8$ by separating the star-forming and quenched populations in the $M_{*}-{\rm SFR}$ plane. In particular, after a detailed inspection of the $M_{*}-{\rm SFR}$ plane, we heuristically flag as star-forming those galaxies with $\log_{10}({\rm SFR}/({\rm M}_{\odot}\,{\rm yr}^{-1}))>0.9\times\log_{10}(M_{*}/{\rm M}_{\odot})-9.2$ at $z\sim 0.365$ (see, Sinigaglia et al., 2022).
@@ 3.2 IllustrisTNG
IllustrisTNG is a suite of cosmological, gravo-magnetohydrodynamical simulations run with the moving-mesh code AREPO (Springel, 2010). IllustrisTNG consists of an improved ad updated version of the original Illustris simulations suite (Nelson et al., 2015), and includes the following baryon physics mechanisms, among others: primordial and metal-line radiative cooling with an ionizing background radiation field, stochastic star formation, an effective equation of state model accounting for the two phases of the interstellar medium, evolution of stellar populations, stellar feedback, seeding and growth of supermassive black holes, AGN feedback (and in particular, energy release from high- and low-accretion rates), and the treatment of magnetic fields.
In this work, we use the following larger-volumes runs: IllustrisTNG-100, run with $N=2\times 1820^{3}$ resolution elements and in volume $V=(75\leavevmode\nobreak\ h^{-1}\leavevmode\nobreak\ {\rm Mpc})^{3}$, and IllustrisTNG-300, run with $N=2\times 2500^{3}$ resolution elements and in volume $V=(205\leavevmode\nobreak\ h^{-1}\leavevmode\nobreak\ {\rm Mpc})^{3}$.
@@ 4 Methods
  • we consider a complete stellar mass sample from the COSMOS2020 stellar mass function at $0.50.2<z<$ 0.5 tabulated in Weaver et al. (2022);
  • we assume the $M_{\rm HI}-M_{\star}$ scaling relation measured via H I spectral stacking from the combination of MIGHTEE and CHILES data sets and presented in Bianchetti et al. (2025). As will be discussed more in detail later on in the paper, this step requires an adequate assumption on the scatter of the $M_{\rm HI}-M_{\star}$ relation. In fact, while the stacking procedure yields the average $M_{\rm HI}$ in each of the studied stellar mass bins, it does not provide any information about the scatter;

### v2

@@ Abstract
Results. We find a deviation of the HIMF at $z\sim 0.37$ from the results at $z\sim 0$ from the ALFALFA survey and at $z\sim 1$ from uGMRT data. Our results for $\Omega_{\rm HI}$ are in broad agreement with other literature results, and follow the overall trend on $\Omega_{\rm HI}$ as a function of redshift. The derived value $\Omega_{\rm HI}=\left(7.02^{+0.59}_{-0.52}\right)\times 10^{-4}$ at $z\sim 0.37$ from the combined analysis deviates at $\sim 2.9\sigma$ from the ALFALFA result at $z\sim 0$.
Conclusions. Our findings about the HIMF and $\Omega_{\rm HI}$ derived from deep, state-of-the-art interferometric surveys differ from previous literature results at $z\sim 0$ and $z\sim 1$, although we are unable to confirm at this stage whether these differences are due to cosmic evolution consistent with a smooth transition of the H I content of galaxies over the last 8 Gyr or due to selection biases and systematics.
@@ 1 Introduction
[…]
@@
In this work, we follow a methodology similar in spirit to the one by Bera et al. (2022) and Chowdhury et al. (2024). In particular, we combine the most recent measurements of the stellar mass function of star-forming galaxies in the COSMOS field at $0.2<z<0.5$ by Weaver et al. (2023) and of the $M_{\rm HI}-M_{\star}$ scaling relation of star-forming galaxies at $0.22<z<0.49$ (median redshift $z\sim 0.37$) by Bianchetti et al. (2025) from the H I stacking analysis of two state-of-the-art deep H I interferometric surveys MIGHTEE (Jarvis et al., 2016) and CHILES (Fernández et al., 2013), both covering the COSMOS field (Scoville et al., 2007). With these tools, we derive novel indirect semi-empirical constraints on the HIMF of star-forming galaxies and the cosmic H I density parameters $\Omega_{\rm HI}$ at $z\sim 0.37$.
@@
Throughout this work, we adopt the following cosmological parameters: $\Omega_{m}=0.3$, $\Omega_{\Lambda}=0.7$, $H_{0}=70.0\penalty 10000\ {\rm km}\penalty 10000\ {\rm s}^{-1}\penalty 10000\ {\rm Mpc}^{-1}$. When relevant, we have rescaled the results from other literature results to the match the cosmological parameters adopted herein.
@@ 2.1 The COSMOS2020 sample and the stellar mass function
The optical catalog underlying the stacking analysis performed by Bianchetti et al. (2025) was built through a cross-match between the COSMOS2020-CLASSIC [^1] Release 1, v2.2, March 2023 photometric catalog (Weaver et al., 2022) and a spectroscopic catalog assembled by merging three different catalogs covering the COSMOS field: the COSMOS spec-$z$ compilation (Khostovan et al., 2025), the DEVILS survey catalog (Davies et al., 2018) and the DESI survey Early Data Release catalog (DESI Collaboration et al., 2024). First, Bianchetti et al. (2025) selected only spectroscopic sources in the COSMOS field (Scoville et al., 2007) and at $0.22<z<0.49$ (median redshift $z\sim 0.37$), i.e. within a subset of the volume probed by the MIGHTEE and CHILES data sets (see Sect. 2.2). In particular, they performed a positional cross-match (with matching radius=$1$ arcsec) between the spectroscopic and the photometric catalog. The photometric catalog includes galaxy properties obtained through spectral energy density (SED) fitting performed with the LePhare code (Ilbert et al., 2006; Arnouts & Ilbert, 2011). Therefore, after the cross-match, they retrieved from the photometric catalog all the aforementioned photometrically-derived parameters for each of the sources in the spectroscopic catalog. Star-forming galaxies were then selected using the in-built color-color selection in the (NUVr)/(rJ)(NUV-r)/(r-J) rest-frame plane. Further quality cuts were then applied to ensure the highest possible reliability, including redshift quality cuts and photometric redshift outliers exclusion.
@@
In this work, we also make use of the results on the stellar mass function in the COSMOS field as measured from the COSMOS2020 catalog (Weaver et al., 2023). In particular, Weaver et al. (2023) performed a detailed study of the stellar mass sample from COSMOS2020, deriving the stellar mass functions for the total, star-forming, quiescent populations at $0.2<z<0.5$, as well as in several higher redshift ranges. The splitting of the galaxy population into star-forming and quiescent was performed according to the same color criterion mentioned above. These aspects, together with the fact that the underlying photometric catalog is the same, make the stellar mass function derived by Weaver et al. (2023) as the ideal candidate to be used in this work. Weaver et al. (2023) used a Markov Chain Monte Carlo method to fit a double Schechter profile to the resulting stellar mass function:
$\begin{split}\Phi\penalty 10000\ d\log_{10}&(M_{\star}/{\rm M}_{\odot})=\log(10)\exp\left(-10^{\log_{10}(M_{\star}/{\rm M}_{\odot})-\log_{10}(\tilde{M}/{\rm M}_{\odot})}\right)\\ &\times\biggl[\tilde{\Phi}_{1}\left(10^{\log_{10}(M_{\star}/{\rm M}_{\odot})-\log_{10}(\tilde{M}/{\rm M}_{\odot})}\right)^{\alpha_{1}+1}\\ &+\tilde{\Phi}_{2}\left(10^{\log_{10}(M_{\star}/{\rm M}_{\odot})-\log_{10}(\tilde{M/{\rm M}_{\odot}})}\right)^{\alpha_{2}+1}\biggr]\penalty 10000\ d\log_{10}(M_{\star}/{\rm M}_{\odot})\quad.\end{split}$ (1)
For the star-forming mass-complete galaxy subsample at $0.2<z<0.5$, Weaver et al. (2023) found the following best-fitting parameters:
$\{\log_{10}(\tilde{M}/{\rm M}_{\odot}),\penalty 10000\ \alpha_{1},\penalty 10000\ \tilde{\Phi}_{1}\times 10^{-3},\penalty 10000\ \alpha_{2},\penalty 10000\ \tilde{\Phi}_{2}\times 10^{-3}\}=\\ =\{10.73^{+0.17}_{-0.15},\penalty 10000\ -1.41^{+0.03}_{-0.04},\penalty 10000\ 0.80^{+1.37}_{-0.56},\penalty 10000\ -0.02^{+0.61}_{-0.79},\penalty 10000\ 0.49^{+0.32}_{-0.33}\}\quad.$
@@ 2.2 The MIGHTEE and CHILES surveys
[…]
CHILES is a deep-field H I survey, carried out with the VLA and imaging H I over a contiguous redshift range of $0<z<0.49$ for the first time (Fernández et al., 2013). The pointing was centered on RA(J2000) 10h01m24s and DEC(J2000) 2d21m00s. It samples a subregion of the COSMOS field; therefore, it provides an independent measurement of the same patch of sky as the MIGHTEE survey. The 25-meter diameter antennas of the VLA provide a field of view (FOV) of around $0.5$ deg at $1.4$ GHz. Due to the rotating VLA configurations, the CHILES observations were split into five observing epochs spaced approximately 15 months apart. In this paper, we use results from successfully processed data from all five of the observing epochs, amounting to approximately $800$ observation hours (approximately $600$ on-source hours). The VLA-B configuration was used, achieving a beam size of approximately $7.2^{\prime\prime}\times 6.4^{\prime\prime}$ at $z\sim 0.37$ ($\approx 38\times 33$ $\rm kpc^{2}$}). The datacube comprises 125 kHz wide spectral channels at $z\sim 0.37$, covering a wide range of frequencies ($950-1420\penalty 10000\ {\rm MHz}$). The RMS ranges between $30$ and $50\penalty 10000\ \mu{\rm Jy\penalty 10000\ beam^{-1}}$ in this range.
@@ 2.3 Summary of stacking results from MIGHTEE and CHILES
$\log_{10}(M_{\star}/{\rm M}_{\odot})$ bin Median $\log_{10}(M_{\star}/{\rm M}_{\odot})$ $N_{\rm gal}$ $\braket{\log_{10}(M_{\rm HI}/{\rm M}_{\odot})}$
@@
$8.0<\log_{10}(M_{\star}/{\rm M_{\odot}})<9.5$ $9.10$ $3507$ $9.51\pm 0.05$
$9.5<\log_{10}(M_{\star}/{\rm M_{\odot}})<9.8$ $9.65$ $909$ $9.68\pm 0.06$
$9.8<\log_{10}(M_{\star}/{\rm M_{\odot}})<10.5$ $10.17$ $1564$ $9.94\pm 0.02$
$\log_{10}(M_{\star}/{\rm M_{\odot}})>10.5$ $10.85$ $636$ $10.07\pm 0.04$
@@
$8.0<\log_{10}(M_{\star}/{\rm M_{\odot}})<9.5$ $9.10$ $2286$ $9.60\pm 0.04$
$9.5<\log_{10}(M_{\star}/{\rm M_{\odot}})<10.5$ $10.06$ $1796$ $9.94\pm 0.02$
$\log_{10}(M_{\star}/{\rm M_{\odot}})>10.5$ $10.85$ $494$ $10.12\pm 0.03$
@@
$8.0<\log_{10}(M_{\star}/{\rm M_{\odot}})<9.5$ $9.10$ $1221$ $9.55\pm 0.07$
$9.5<\log_{10}(M_{\star}/{\rm M_{\odot}})<10.5$ $10.03$ $767$ $9.82\pm 0.06$
$\log_{10}(M_{\star}/{\rm M_{\odot}})>10.5$ $10.83$ $142$ $10.13\pm 0.08$
Table 1: Data points from the H I spectral stacking analysis performed by Bianchetti et al. (2025). The first column reports the $M_{\star}$ interval underlying each of the measured data point, the second column the corresponding median $M_{\star}$, the third column the number of stacked galaxies, and the fourth column the average $M_{\rm HI}$ measured from stacking.
@@ 3.1 SIMBA
In this work, we make use of the publicly-available galaxy catalogs at $z\sim 0.365$ for SIMBA-100 and for all the SIMBA-50 boxes. Consistently with the observational sample selection by Bianchetti et al. (2025), we select only star-forming galaxies with $\log_{10}(M_{*}/{\rm M}_{\odot})>8$ by separating the star-forming and quenched populations in the $M_{*}-{\rm SFR}$ plane. In particular, after a detailed inspection of the $M_{*}-{\rm SFR}$ plane, we heuristically flag as star-forming those galaxies with $\log_{10}({\rm SFR}/({\rm M}_{\odot}\,{\rm yr}^{-1}))>0.9\times\log_{10}(M_{*}/{\rm M}_{\odot})-9.2$ at $z\sim 0.365$ (see, Sinigaglia et al., 2022).
@@ 3.2 IllustrisTNG
IllustrisTNG is a suite of cosmological, gravo-magnetohydrodynamical simulations run with the moving-mesh code AREPO (Springel, 2010). IllustrisTNG consists of an improved and updated version of the original Illustris simulations suite (Nelson et al., 2015), and includes the following baryon physics mechanisms, among others: primordial and metal-line radiative cooling with an ionizing background radiation field, stochastic star formation, an effective equation of state model accounting for the two phases of the interstellar medium, evolution of stellar populations, stellar feedback, seeding and growth of supermassive black holes, AGN feedback (and in particular, energy release from high- and low-accretion rates), and the treatment of magnetic fields.
In this work, we use the following larger-volumes runs: IllustrisTNG-100, run with $N=2\times 1820^{3}$ resolution elements and in volume $V=(75\penalty 10000\ h^{-1}\penalty 10000\ {\rm Mpc})^{3}$, and IllustrisTNG-300, run with $N=2\times 2500^{3}$ resolution elements and in volume $V=(205\penalty 10000\ h^{-1}\penalty 10000\ {\rm Mpc})^{3}$.
@@ 4 Methods
  • we consider a complete stellar mass sample from the COSMOS2020 stellar mass function at $0.2<z<0.5$ tabulated in Weaver et al. (2022);
  • we assume the $M_{\rm HI}-M_{\star}$ scaling relation measured via H I spectral stacking from the combination of MIGHTEE and CHILES data sets and presented in Bianchetti et al. (2025). As will be discussed in more detail later on in the paper, this step requires an adequate assumption on the scatter of the $M_{\rm HI}-M_{\star}$ relation. In fact, while the stacking procedure yields the average $M_{\rm HI}$ in each of the studied stellar mass bins, it does not provide any information about the scatter;